OpenStudy (anonymous):

What is the solution set of |x + 3| – 2 = –5?

OpenStudy (jamesj):

You see that |x+3| = -3 Do you agree so far?

OpenStudy (anonymous):

I have no idea what to do, honestly.

OpenStudy (jamesj):

Ok. So, I added 2 to each side |x+3| - 2 + 2 = -5 + 2 Hence |x+3| + 0 = -3 |x+3| = -3 Yes?

OpenStudy (anonymous):

Oh. okay. I see now...

OpenStudy (anonymous):

either x+3 or -x-3 would be used to replace the absolute value sign so.. x + 3 - 2 = 5 or -x - 3 - 2 = 5 solve

OpenStudy (jamesj):

Good. Now we have a problem.

OpenStudy (jamesj):

@gspade: totally wrong, sorry. Is there ANY number y such that |y| = -3 ?

OpenStudy (jamesj):

Because |3| = 3 and |-3| = 3. So is there any number y such that |y| = -3 ?

OpenStudy (anonymous):

agree with JamesJ

OpenStudy (jamesj):

So to recap. So far we shown the original problem is equivalent to the problem of find x such |x+3| = -3 Now the question I am raising is this: can that equation even have any solutions in theory?

OpenStudy (jamesj):

That is, are there ANY numbers y that have the property that | y | = -3 ?

OpenStudy (jamesj):

Ok. You've buggered off. But think about what I'm asking. Good luck.